OpenStudy (anonymous):

The sum of two consecutive odd integers is 236. What is the smaller integer?

OpenStudy (kainui):

Well if you have consecutive odd numbers then they must be two apart. For example 5 and 7 are consecutive odd numbers right, so if we add 2 to 5 we get 7. The next consecutive odd number is 9 right, and 7+2=9. So now we can apply this rule to a variable we don't know yet. We don't know what the number x is, but we're saying that if it's odd then the next odd number is x+2. Now using (x) and (x+2) since these are our consecutive odd numbers, can you put these into some kind of relationship to 236 based on the information given?

OpenStudy (anonymous):

So would you divide the 2 into the 236?

OpenStudy (muzzack):

let x be the smaller integer x + (x + 2) = 236 2x + 2 = 236 2x = 236 - 2 2x = 234 x = 117 => smaller integer

OpenStudy (anonymous):

Ooh. I Understand Know. Thanks To Both ! (: